How to solve systems of linear equations such as this one.
4x - y = -4
12 - 3y = 24

Answer by math-vortex(648)

How to solve systems of linear equations such as this one.
4x - y = -4
12 - 3y = 24
==========================
I wonder if you meant 12x - 3y = 24 rather than 12 - 3y = 24.
.
I will solve the problem you typed in. If you need more help, post a new question, or send me a message.
.
You have two linear equations here. On a coordinate graph, the are two straight lines which intersect a single point (x, y). We will use algebra to find the coordinates of the intersection point.
.
First, let's simplify the second equation:
12 - 3y = 24
.
Subtract 12 from both sides.
-3y = 24 - 12
-3y = 12
.
Divide both sides by -3.
y = 12/(-3)
y = -4
.
Now we can substitute -4 for y in the first equation, and solve for x.
4x - y = -4
4x - (-4) = -4
4x + 4 = -4
4x = -8
x = -2
.
Since x = -2 and y = -4 our solution is the ordered pair (-2,-4).
.
We should check our work with the original equations. The ordered pair (-2,-4) will make both equations true.
.
4x - y = -4
4(-2) - (-4) = -4
-8 + 4 = -4
-4 = -4 (true!)
.
12 - 3y = 24
12 - 3(-4) = 24
12 - (-12) = 24
12 + 12 = 24
24 = 24 (true!)
.
Hope this helps!
.
Ms.Figgy
math.in.the.vortex